OpenStudy (anonymous):

Show that for all integers m and n, with m ≠ +/-n, the integral from -π to π of cos(mθ)sin(nθ) dθ = 0

OpenStudy (anonymous):

\[\cos p \sin q=\frac{1}{2}(\sin(p+q)-\sin(p-q))\]

OpenStudy (anonymous):

use that and then integrate

OpenStudy (anonymous):

so 1/2(sin(mθ+nθ)-sin(mθ+nθ)) derived would be (1/2)(cos(mθ+nθ)(mθdθ)(nθdθ)-cos(mθ+nθ)(mθdθ)(nθdθ)) is that right?

OpenStudy (anonymous):

and then plug in π for θ minus the quantity of -π for θ

OpenStudy (anonymous):

emm....u must integrate :/

OpenStudy (anonymous):

hah oops

OpenStudy (anonymous):

=\[\int\limits_{-π}^{π}sin(mθ+nθ)]-\int\limits_{-π}^{π}sin(mθ-nθ)\]

OpenStudy (anonymous):

yes integrate and apply bounds

OpenStudy (anonymous):

one another point...integrand is an odd function and ur interval is symmetric with respect to y-axis

OpenStudy (anonymous):

i honestly don't even know where to begin integrating those. i suck at trig integrals

OpenStudy (anonymous):

nevermind, i figured it out, thanks!